OpenStudy (anonymous):

find the equation of a line that goes through points (2,4) and (6,8)

OpenStudy (anonymous):

To find the equation of a line you need a point it passes through and it's gradient. to get the gradient you use the formula \[\frac{ y_{2}-y_{1}}{ x_{2}-x_{1} }\] Which comes out to be\[\frac{ 8-4 }{ 6-2 } = \frac{ 4 }{ 4 } = 1\] Now we know the gradient you can put it all in the formula for any general line\[y-y_{1}=m(x-x _{1})\] So using the point 2,4, \[y-4=1(x-2) -> y-4=x-2 -> y=x-2+4 -> y=x+2\] Which is the answer in y=mx+c form :D

OpenStudy (anonymous):

 can you find the slope of this line?

OpenStudy (anonymous):

That is not my question

OpenStudy (anonymous):

Can somebody please answer what are the 3 possible points you can use to graph y= 3-x

OpenStudy (anonymous):

When x = 0 is the first point so y = 3-0 y = 3. So the first point is (0,3) Second is when y = 0 0= 3-x, x=3 so the second is (3,0) Because the line is a straight line all you need is two points and draw a line through them with a ruler. If you really need a third point just read one off the line :)